OpenStudy (sen_turtle):

you are building a rectangular brick patio surrounded by crushed stone in a rectangular courtyard as shown. The crushed stone boarder has a uniform width x (in feet). You have enough money in your budget to purchase patio bricks to cover 140 sq. feet. Solve the equation 140=(20-2x) (16-2x) to find the width of the border.

OpenStudy (paxpolaris):

\[ 140=(20-2x)(16-2x) \] Divide both sides by 4:\[ {140\over 4}={(20-2x)\over 2}\times{(16-2x) \over 2}\]\[35=\left( 10-x \right)\left( 8-x \right)\]

OpenStudy (paxpolaris):

\[35 = 80-18x+x^2\]

OpenStudy (paxpolaris):

\[x^2-18x+45=0\]

OpenStudy (paxpolaris):

solve by factoring to get 2 answers ... only one works

OpenStudy (paxpolaris):



OpenStudy (sen_turtle):

Okay, thank you!